# Identifying Proca-star mergers via consistent ultralight-boson mass estimates across gravitational-wave events

### v1

@@ V Discussion
Evidence for a given type of GW source can be obtained either through strong evidence coming from a single a observation or through the accumulation of mild evidence across different events. GW sources like boson stars or neutron stars are described by characteristic parameters, respectively the underlying boson mass and the tidal deformability, which shall common across the whole source population.
@@
Focusing on the case of a particular class of boson stars known as Proca stars, we have shown how boson-mass consistencies across events can be exploited to detect underlying populations of Proca stars and their properties even if no individual event can be conclusively identified as such. In particular, our formalism allows to progressively constrain the number of boson-star families and their characteristic masses as observations accumulate. The same formalism could be easily adapted to, for instance, constrain the number of families of compact objects populating the Universe characterised by e.g. their characteristic tidal deformabilities or the number and type of possible equations of state characterising neutron-stars and alternative compact objects.
@@
Applied to the set of events we analyse, we conclude that, these are most consistent with all being black-hole mergers, i.e., we do not need to invoke any number of ultralight bosons to explain then. We note, however, that a number $N_{\rm bosons}=1$ and $N_{\rm bosons}=2$ ultralight bosons are not conclusively discarded. Moreover, we show that due to the consistency of the boson-mass estimates across these events, a value of $N_{\rm bosons}=1$ would be conclusively preferred after 5 to 9 observations of similar events. We also show that such evidence would not be achieved if ignoring the mentioned mass consistencies, which underscores the power of our formalism.
@@
Consistently with previous work, we have checked that our results result are influenced by the intrinsic weakness of the Proca-star merger configurations with which the events have been compared, which imposes a heavy penalty on the corresponding evidence. Removing this effect, we see that the hypothesis that all the studied events are black hole mergers (i.e. $N_{\rm bosons}=0$) is mildly rejected with Bayes Factors of $5$($40$) when the event S200114 is considered (discarded) as a true GW signal.
@@
We note that our results are heavily by the limited number of currently available Proca-star merger simulations and, even more importantly, the restriction of these to almost “toy models” restricted to head-on mergers First, such configurations are astrophysically very unlikely. Second, as mentioned before, these lead to very weak signals that require the source to be way closer than a black-hole merger producing the same signal, which imposes a heavy statistical penalty. Third, the shortness of the signals limits our study to very short GW signals from the heaviest black-hole merger detections, preventing the analysis of lighter systems with visible inspirals. Current work e.g. Palloni et al. (2025) is ongoing towards simulating less eccentric configurations. Finally, the 769 numerical simulations to which our events were compared, omit the impact of varying the relative phase of the fields the stars at merger for fixed configurations, which is determined by the relative phase at the start of the simulation. While in the simulations used in our input studies the relative phase of the stars is always fixed to zero, Ref. Sanchis-Gual et al. (2022) showed that varying this can have a dramatic impact in the signal. The future development of extended catalogs of numerical simulations and continuous surrogate models Luna et al. (2024) together with the increased sensitivity of the existing gravitational-wave detector network and the advent of next generation detectors such as LISA, Cosmic Explorer and Einstein Telescope Amaro-Seoane et al. (2017); Reitze et al. (2019); Hild et al. (2011); Abac et al. (2025) shall reveal or rule out the existence of compact exotic objects and, in particular, that of Proca Stars.
@@ Acknowledgements
ALM is supported by the Spanish Agencia Estatal de Investigación through the grant PRE2022-102569, funded by MCIN/AEI/10.13039/ 501100011033 and the FSE+. This work has received financial support from María de Maeztu grant CEX2023-001318-M funded by MICIU/AEI/10.13039/501100011033, and from the Xunta de Galicia (CIGUS Network of Research Centres) and the European Union. JCB is supported by a the a Ramón y Cajal Fellowship RYC2022-036203-I from the Spain-Ministerio de Ciencia e Innovación. JCB acknowledges further support by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. SHWL acknowledges support by grants from the Research Grants Council of Hong Kong (Project No. CUHK 14304622 and 14307923), and the Direct Grant for Research from the Research Committee of The Chinese University of Hong Kong. The authors acknowledge computational resources provided by the CIT cluster of the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY0823459; and the support of the NSF CIT cluster for the provision of computational resources for our parameter inference runs. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This manuscript has LIGO DCC number P2500136.

### v2

@@ V Discussion
Evidence for a given type of GW source can be obtained either through strong evidence coming from a single a observation or through the accumulation of mild evidence across different events. GW sources like boson stars or neutron stars are described by characteristic parameters, respectively the underlying boson mass and the tidal deformability, which shall have common value across the whole source population.
@@
Focusing on the case of a particular class of boson stars known as Proca stars, we have shown how parameter consistencies across events ( in our case, the boson masss) can be exploited to detect underlying populations of Proca stars and their properties even if no individual event can be conclusively identified as such. In particular, our formalism allows to progressively constrain the number of boson-star families and their characteristic masses as observations accumulate. The same formalism could be easily adapted to, for instance, constrain the number of families of compact objects populating the Universe characterised by e.g. their characteristic tidal deformabilities or the number and type of possible equations of state characterising neutron-stars and alternative compact objects.
@@
Applied to the set of events we analyse, we conclude that, these are most consistent with all being black-hole mergers, i.e. we do not need to invoke any number of ultralight bosons to explain then. We note, however, that a number $n=1$ and $n=2$ ultralight bosons are not conclusively discarded. Moreover, we show that due to the consistency of the boson-mass estimates across these events, our framework would make a value of $n=1$ be conclusively preferred after 5 to 9 observations of events with similar characteristics. We also show that such evidence would not be achieved if ignoring the mentioned mass consistencies, which underscores the power of our formalism.
@@
Consistently with previous work, we have checked that our results are heavily influenced by the intrinsic weakness of the Proca-star merger configurations with which the GW events have been compared, which imposes a strong penalty on the model. The impact of this ingredient is reflected in the fact that, if removing it, the hypothesis that all the studied events are black hole mergers (i.e. $n=0$) is mildly rejected with Bayes Factors of 55(4040) when the event S200114 is considered (discarded) as a true GW signal.
@@
We note that our results are also heavily driven by the limited number of currently available Proca-star merger simulations and, even more importantly, by the restriction of these simulations to almost “toy models” restricted to head-on mergers . First, such configurations are astrophysically very unlikely. Second, as mentioned before, these lead to very weak signals that require the source to be way closer than a black-hole merger producing the same signal, which imposes a heavy statistical penalty. Third, the extremely short duration of the these signals limits our study to very short GW signals from the heaviest black-hole merger detections, preventing the analysis of lighter systems with visible inspirals. Current work, e.g. Ref. Palloni et al. (2025), is ongoing towards simulating less eccentric configurations. Finally, the 769 numerical simulations to which our events were compared omit the impact of varying the relative phase of the fields of the stars at merger. While in the simulations used in our input studies the relative phase of the stars is always fixed to zero, Ref. Sanchis-Gual et al. (2022) showed that varying this can have a dramatic impact in the signal. The future development of extended catalogs of numerical simulations and continuous surrogate models Luna et al. (2024), together with the increased sensitivity of the existing gravitational-wave detector network and the advent of next generation detectors such as LISA, Cosmic Explorer and Einstein Telescope Amaro-Seoane et al. (2017); Reitze et al. (2019); Hild et al. (2011); Abac et al. (2025) shall reveal or rule out the existence of compact exotic objects and, in particular, that of Proca Stars.
@@ Acknowledgements
ALM is supported by the Spanish Agencia Estatal de Investigación through the grant PRE2022-102569, funded by MCIN/AEI/10.13039/ 501100011033 and the FSE+. This work has received financial support from María de Maeztu grant CEX2023-001318-M funded by MICIU/AEI/10.13039/501100011033, and from the Xunta de Galicia (CIGUS Network of Research Centres) and the European Union. JCB is supported by the Ramon y Cajal Fellowship RYC2022-036203-I, by the Grant PID2024-160643NB-I00 of the Spanish Ministry of Science, Innovation and Universities; and by the Grant ED431F 2025/04 of the Galician Conselleria de Educación, Ciencia, Universidades e Formación Profesional. IGFAE is supported by the Ayuda Maria de Maeztu CEX2023-001318-M funded by MICIU/AEI /10.13039/501100011033. We acknowledge further support from the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. SHWL acknowledges support by grants from the Research Grants Council of Hong Kong (Project No. CUHK 14304622 and 14307923), and the Direct Grant for Research from the Research Committee of The Chinese University of Hong Kong. The authors acknowledge computational resources provided by the CIT cluster of the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY0823459; and the support of the NSF CIT cluster for the provision of computational resources for our parameter inference runs. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This manuscript has LIGO DCC number P2500136.